OpenStudy (anonymous):

Write the solution to the following problem as an ordered pair: The sum of two numbers is 170 and their difference is 26.

OpenStudy (anonymous):

assume that the pair of numbers are x and y x+y= 170 x-y= 26 x= 26+y 26+y+y=170 26+2y=170 2y=170-26 2y=144 y=72 x=26+y x=26+72 x=98 the numbers are 98 and 72 to check if the answers is right 98+72=170 98-72=26